# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ Abstract.
For an exact twist map $f$, we introduce an inherent Lipschitz dynamics $\Sigma_{+}$ given by the discrete forward Lax-Oleinik semigroup. We investigate several properties of $\Sigma_{+}$ and show that for any discrete weak K.A.M. solution $u$, the non-differentiable points of $u$ are globally propagated and forward invariant by $\Sigma_{+}$. In particular, such propagating dynamics possesses the same rotation number as the associated Aubry-Mather set with respect to $u$.
A detailed exposition of the corresponding Arnaud’s observation [Arn11] is then provided via $\Sigma_{+}$. Furthermore, we construct and analyze the dynamics on the pseudo-graphs of discrete weak K.A.M. solutions.
@@ 1. Introduction
The celebrated Aubry-Mather theory developed independently by S. Aubry [Aub83, ALD83] and J. Mather [Mat82] provides a novel approach to the study of the dynamics of twist diffeomorphisms of the annulus. The existence of action minimizing sets (referred as Aubry-Mather sets) –which could be invariant circles or invariant Cantor sets etc–is established. Moreover, one may have
@@
  • all the Aubry-Mather sets are vertically ordered in the annulus according to the rotation numbers. See [Gol01].
@@
Let $c\in\mathbb{R}$ be some cohomology class [^1] In the present paper, the cohomology class $c$ is fixed at the beginning and one can assume $c=0$ for simplicity. The reason why we insist writing $c$ explicitly in the introduction is that we will apply the results here for future work. of $\mathbb{T}$. One can also define the lift $F_{c}$ and the corresponding generating function $S_{c}$ at cohomology class $c$ given by
@@
  • Mather’s $\alpha$-function: $H^{1}(\mathbb{T},\mathbb{R})\rightarrow\mathbb{R},c\mapsto-\bar{S}_{c}$, which is differentiable everywhere.
@@
Therefore, we could state our first result below:
@@
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the map $\Sigma_{+}:\mathbb{R}\rightarrow\mathbb{R},\ x\mapsto y$ is well-defined and implicitly given by
$T^{+}[u_{c}](x)=u_{c}(y)-S_{c}(x,y).$
Moreover, the map $\Sigma_{+}$ satisfies the following properties:
@@
As an application of the above properties of the singular dynamics, we immediately obtain that the singularities of a discrete weak K.A.M. solution $u$ can be globally propagated with a rotation number.
@@
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Let $x_{0}\in\operatorname{Sing}(u_{c})$ and one can define the one-sided sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ iteratively by:
$x_{n+1}:=\mathop{\arg\max}_{y\in\mathbb{R}}\{u_{c}(y)-S_{c}(x_{n},y)\}\qquad\text{ for any }n\in\mathbb{Z}_{+}.$
@@
If $\operatorname{Sing}(u_{c})\neq\emptyset$, we can pick $x_{0}\in\operatorname{Sing}(u_{c})$ and define $x_{n+1}:=\Sigma_{+}(x_{n}),\ \forall n\in\mathbb{Z}_{+}$. In addition, if $\operatorname{Sing}(u_{c})/\mathbb{Z}$ is a finite set, we know that there exist two different natural numbers $n<m$ such that $x_{m}-x_{n}\in\mathbb{Z}$. Then we obtain that $\rho_{c}=(x_{m}-x_{n})/(m-n)\in\mathbb{Q}$. Hence $\alpha^{\prime}(c)$ also belongs to $\mathbb{Q}$.
@@
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
@@
Following Arnaud and Zavidovique’s article [AZ23], we call $\operatorname{Graph}(c+\nabla^{+}u_{c})$ the full pseudo-graph of a weak K.A.M. solution $u_{c}$ at cohomology class $c$. If $x\in\operatorname{Sing}(u_{c})$, then $\nabla^{+}u_{c}(x)$ is not a singleton and we refer to $\{x\}\times(c+\nabla^{+}u_{c}(x))$ as a vertical segment of the full pseudo-graph $\operatorname{Graph}(c+\nabla^{+}u_{c})$. Notice that there is a natural forward map
@@
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
Figure 2. Singular dynamics on the full pseudo-graph $\operatorname{Graph}(c+\nabla^{+}u_{c})$ of a discrete weak K.A.M. solution $u_{c}$ at cohomology class $c$, where different colors are used to represent the corresponding relationships before and after the map $\Lambda_{c}$.
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
Figure 3. Singular dynamics on the $\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$, where different colors are used to represent the corresponding relationships before and after the map $\Omega_{c}$.
@@
    $\Lambda_{c}$ does not reverse the orientation of the full pseudo-graph.
@@
Observing Figure 2 reveals that within the singular dynamics, certain points are attracted to the singularities. However, it is also evident that some points do not propagate into the singularities during the iteration of $\Lambda_{c}$. Thus, one could conclude:
@@
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the two sets
$c+\mathcal{I}^{*}(u_{c}):=\bigcap_{n\geq 0}F^{-n}(\overline{\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u_{c})});$
$c+\mathcal{K}^{*}(u_{c}):=\bigcup_{\begin{subarray}{c}x\in\operatorname{Sing}(u_{c})\\ p\in c+\nabla^{+}u_{c}(x)\end{subarray}}\Lambda_{c}(x,p)$
satisfy the following properties
  1. (i)
    the set $c+\mathcal{I}^{*}(u_{c})$ is closed, non-empty and $\Lambda_{c}$-invariant;
  2. (ii)
    the restriction of $\Lambda_{c}$ on $c+\mathcal{I}^{*}(u_{c})$ is the lift $F$ of the twist map $f$;
  3. (iii)
    the set $c+\mathcal{K}^{*}(u_{c})$ is at most countable and contained within the vertical segments of $\operatorname{Graph}(c+\nabla^{+}u_{c})$;
  4. (iv)
    for any $(x,p)\in\operatorname{Graph}(c+\nabla^{+}u_{c})\setminus(c+\mathcal{I}^{*}(u_{c}))$, there exists $n\in\mathbb{Z}_{+}$ such that
    $\Lambda_{c}^{n}(x,p)\in c+\mathcal{K}^{*}(u_{c}).$
Before ending this section, as an application of $c+\mathcal{I}^{*}(u_{c})$, we represent the Aubry set in Theorem 2.15 that:
$c+\mathcal{A}^{*}_{c}=\bigcap\left(c+\mathcal{I}^{*}(u_{c})\right),$
where the intersection is considered over all discrete weak K.A.M. solutions $u_{c}$ at the cohomology class $c$. This constitutes the weak Hamilton-Jacobi theorem (see e.g. [Fat08]) in the discrete setting. For the sake of simplicity, all proofs in Sections 3-5 only deals with the case where $c=0$.
2. Preliminary
@@
  • For the product space $X\times Y$, the maps $\pi_{1}:X\times Y\rightarrow X$ and $\pi_{2}:X\times Y\rightarrow Y$ are the projections on the first and second variable.
@@
2.1. Discrete weak K.A.M. theory on $d$-dimensional torus
Let us recall some fundamental definitions and properties on semiconcave functions (see [CS04, Chapter 3] for instance).
Definition 2.1.
Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function.
  • For any $x\in\mathbb{R}^{d}$, the sets
  • Let $\Omega\subset\mathbb{R}^{d}$ be a convex open set, we say that $u$ is semiconcave (with a linear modulus) on $\Omega$ if there exists a constant $C\geq 0$ such that
[…]
  • For any $x\in\mathbb{R}^{d}$, a covector $p\in(\mathbb{R}^{d})^{*}$ is called a reachable differential of $u$ at $x$ if a sequence $\{x_{n}\}\subset\mathbb{R}^{d}\backslash\{x\}$ exists such that $u$ is differentiable at $x_{k}$ for each $k\in\mathbb{Z}_{+}$, and
    $\lim_{k\rightarrow\infty}x_{k}=x,\quad\lim_{k\rightarrow\infty}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u\left(x_{k}\right)=p.$
    The set of all reachable gradients of $u$ at $x$ is denoted by *$\nabla^{*}u(x)$.
Proposition 2.2.
  1. (i)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be semiconcave function on $\Omega$ with linear modulus and constant $C$. Then
    $u(x)-u(x_{0})\leq p_{0}(x-x_{0})+\frac{C}{2}|x-x_{0}|^{2}$
    for any $x,x_{0}\in\Omega$ and $p_{0}\in\nabla^{+}u(x_{0})$.
  2. (ii)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function, $\Omega\subset\mathbb{R}^{d}$ be a convex open set and $x_{0}\in\Omega$. If there exists $C\geq 0$ and $p_{0}\in(\mathbb{R}^{d})^{*}$ such that
    $u(x)-u(x_{0})\leq p_{0}(x-x_{0})+\frac{C}{2}|x-x_{0}|^{2},\ \quad\forall x\in\Omega$
    then $p_{0}\in\nabla^{+}u(x_{0})$.
  3. (iii)
    Let $u,v:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be two continuous functions, $\Omega\subset\mathbb{R}^{d}$ be a convex open set and $x\in\Omega$. If $u\leq v$ on $\Omega$ and $u(x)=v(x)$, then $\nabla^{+}u(x)\supset\nabla^{+}v(x)$ and $\nabla^{-}u(x)\subset\nabla^{-}v(x)$.
  4. (iv)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function and $x\in\mathbb{R}^{d}$. Then $u$ is differentiable at $x$ if and only if $\nabla^{+}u(x)$ and $\nabla^{-}u(x)$ are both nonempty; in this case we have that
  5. (v)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be semiconcave function on $\Omega$. Then *$\nabla^{*}u(x)\subset\partial\nabla^{+}u(x)$ for all $x\in\Omega$, where $\partial$ is the boundary taken with respect to the standard topology of $\mathbb{R}^{d}$.
@@
    $S$ is diagonal-periodic:
  5. (5)
    $S$ is ferromagnetic; that is, the two maps $x\mapsto\partial_{2}S(x,y)$ and $y\mapsto\partial_{1}S(x,y)$ are homeomorphisms between $\mathbb{R}^{d}$ and $(\mathbb{R}^{d})^{*}$ for all $(x,y)$.
@@
The notion of ferromagnetic terminology is introduced by Garibaldi and Thieullen (see [GT11, Definition 2.4]). If $S$ satisfies Hypothesis 1, then the discrete standard map $F:\mathbb{R}^{d}\times(\mathbb{R}^{d})^{*}\righttoleftarrow$ can be defined by
$F(x,p)=(y,p^{\prime}):\Longleftrightarrow\begin{cases}p=-\partial_{1}S(x,y)\\ p^{\prime}=\partial_{2}S(x,y)\end{cases}.$
@@
We recall the discrete backward Lax-Oleinik semigroup and discrete forward Lax-Oleinik semigroup defined in the introduction:
$\forall x\in\mathbb{R}^{d},\quad T^{-}[u](x):=$ $\inf_{y\in\mathbb{R}^{d}}\big{\{}u(y)+S(y,x)\big{\}},$

### v2

@@ Abstract.
For any exact twist map $f$ and any cohomology class $c\in\mathbb{R}$, let $u_{c}$ be any associated discrete weak K.A.M. solution, and we introduce an inherent Lipschitz dynamics $\Sigma_{+}$ given by the discrete forward Lax-Oleinik semigroup. We investigate several properties of $\Sigma_{+}$ and show that the non-differentiable points of $u_{c}$ are globally propagated and forward invariant by $\Sigma_{+}$. In particular, such propagating dynamics possesses the same rotation number $\alpha^{\prime}(c)$ as the associated Aubry-Mather set at cohomology class $c$.
As applications, we provide via $\Sigma_{+}$ a discrete analogue of Bernard’s regularization theorem [Ber07] and a detailed exposition of Arnaud’s observation [Arn11]. Furthermore, we construct and analyze the corresponding dynamics on the full pseudo-graphs of discrete weak K.A.M. solutions.
@@
The datasets generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
@@ 1. Introduction
In the 1890 volume of Acta Mathematica, H. Poincaré introduce some of the basic ideas about twist maps of the annulus while he researched on the stability of orbits of the three body problem.
We focus on some recent developments on the twist maps. The celebrated Aubry-Mather theory developed independently by S. Aubry [Aub83, ALD83] and J. Mather [Mat82] provides a novel approach to the study of the dynamics of twist diffeomorphisms of the annulus (a weak analogue for twist homeomorphisms can be found in [KO97]). Especially, the existence of action minimizing sets (referred as Aubry-Mather sets) –which could be invariant circles or invariant Cantor sets etc–is established. Moreover, one may have
@@
  • all the Aubry-Mather sets are vertically ordered in the annulus according to the rotation numbers. See e.g. [Gol01].
@@
Let $c\in\mathbb{R}$ be some cohomology class [^1] In the present paper, the cohomology class $c$ is fixed at the beginning and one can assume $c=0$ for simplicity. The reason why we insist writing $c$ explicitly in the introduction is that we will apply the results here for future work, such as establishing $c$-equivalence tools (see [CY04]) for searching for diffusion orbits in the discrete setting. of $\mathbb{T}$. One can also define the lift $F_{c}$ and the corresponding generating function $S_{c}$ at cohomology class $c$ given by
@@
  • Mather’s $\alpha$-function: $H^{1}(\mathbb{T},\mathbb{R})\rightarrow\mathbb{R},c\mapsto-\overline{S}_{c}$, which is differentiable everywhere and so the corresponding rotation number equals $\alpha^{\prime}(c)$.
@@
We now state our first result below:
@@
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the optimal forward map
$\Sigma_{+}:\mathbb{R}\rightarrow\mathbb{R},\ x\mapsto\mathop{\arg\max}_{y\in\mathbb{R}}\{u_{c}(y)-S_{c}(x_{n},y)\}$
is single-valued. Moreover, the map $\Sigma_{+}$ satisfies the following properties:
@@
As an application of the above properties of the singular dynamics, we immediately obtain that the singularities of any discrete weak K.A.M. solution can be globally propagated with a rotation number, which answers a question of Wei Cheng [Che21].
@@
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Let $x_{0}\in\operatorname{Sing}(u_{c})$ and one can define the one-sided sequence $\{x_{n}\}_{n\in\mathbb{N}}$ iteratively by:
$x_{n+1}:=\mathop{\arg\max}_{y\in\mathbb{R}}\{u_{c}(y)-S_{c}(x_{n},y)\}\qquad\text{ for any }n\in\mathbb{N}.$
@@
If $\operatorname{Sing}(u_{c})\neq\emptyset$, one can pick $x_{0}\in\operatorname{Sing}(u_{c})$ and define iteratively $x_{n+1}:=\Sigma_{+}(x_{n}),\ \forall n\in\mathbb{N}$. In addition, if $\operatorname{Sing}(u_{c})/\mathbb{Z}$ is a finite set, we know that there exist two different natural numbers $n<m$ such that $x_{m}-x_{n}\in\mathbb{Z}$. Then we obtain that $\rho_{c}=(x_{m}-x_{n})/(m-n)\in\mathbb{Q}$. Hence $\alpha^{\prime}(c)$ also belongs to $\mathbb{Q}$.
@@
We can use the singular dynamics $\Sigma_{+}$ to present a discrete analogue of Patrick Bernard’s regularization theorem [Ber07]. This extends the results of [Zav23].
Theorem 1.4.
For any bounded function $u:\mathbb{T}\rightarrow\mathbb{R}$, the function $T^{+}\circ T^{-2}[u]$ is $C^{1,1}$.
@@
$\Sigma_{+}$$c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}]$$c+\nabla^{+}u_{c}$
@@
Following [AZ23], we call $\operatorname{Graph}(c+\nabla^{+}u_{c})$ the full pseudo-graph of a weak K.A.M. solution $u_{c}$ at cohomology class $c$. If $x\in\operatorname{Sing}(u_{c})$, then $\nabla^{+}u_{c}(x)$ is not a singleton and we refer to $\{x\}\times(c+\nabla^{+}u_{c}(x))$ as a vertical segment of the full pseudo-graph $\operatorname{Graph}(c+\nabla^{+}u_{c})$. Notice that there is a natural forward map
@@
$y$$x_{-}(y)$$x_{+}(y)$$\Sigma_{+}(y)$$\operatorname{Graph(}c+\nabla^{+}u_{c})$$x_{-}(\Sigma_{+}(y))$$x_{+}(\Sigma_{+}(y))$$\Lambda_{c}$$\Lambda_{c}$
Figure 2. Singular dynamics on the full pseudo-graph $\operatorname{Graph}(c+\nabla^{+}u_{c})$ of a discrete weak K.A.M. solution $u_{c}$ at cohomology class $c$, where different colors are used to represent the corresponding pieces before and after the map $\Lambda_{c}$.
$\operatorname{Graph(}c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$$\Omega_{c}$$y$$x_{-}(y)$$x_{+}(y)$$(\Sigma_{+})^{-2}(y)$$\Omega_{c}$
Figure 3. Singular dynamics on the $\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$, where different colors are used to represent the corresponding pieces before and after the map $\Omega_{c}$.
@@
    $\Lambda_{c}$ preserves the orientation of the full pseudo-graph.
@@
Heuristically, from Figure 2, one could observe that some points are attracted to the singularities of $u_{c}$ under the singular dynamics $\Lambda_{c}$. However, on the other hand, it is also evident that some other points do not propagate into the singularities during the iteration of $\Lambda_{c}$. Thus, one could conclude:
@@
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. For any discrete weak K.A.M. solution $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ for $f$ at cohomology class $c$, the $\alpha$-limit set of $\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u_{c})$ under $\Lambda_{c}$:
$\alpha(\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u_{c})):=\bigcap_{n\in\mathbb{Z}_{+}}\overline{\{\Lambda^{i}_{c}(x,p)\mid i\leq-n,(x,p)\in\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u_{c})\}}$
is closed, non-empty and invariant under the dynamics $F$. Furthermore, the Aubry set at cohomology class $c$ can be expressed as
$\mathcal{A}^{*}_{c}=\bigcap_{u_{c}}\alpha(\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u_{c}))-(0,c),$
where the intersection is taken over all discrete weak K.A.M. solutions $u_{c}$ at cohomology class $c$.
We can extend some classical results [Kat95, Proposition 11.2.5] for orientation-preserving circle homeomorphism to orientation-preserving continuous circle map of degree $1$, without requiring invertibility. Note that $\Sigma_{+}$ is precisely such a map. Consequently, if $\alpha^{\prime}(c)\in\mathbb{R}\setminus\mathbb{Q}$, the $\alpha$-limit set under $\Sigma_{+}$ is either $\mathbb{R}$ or a Cantor set. The same conclusion holds for the $\alpha$-limit set under $\Lambda_{c}$, since $\Sigma_{+}\circ\pi_{1}=\pi_{1}\circ\Lambda_{c}$. Furthermore, a theorem due to Mather [Mat90] and Bangert [Ban88] guarantees the uniqueness of discrete weak K.A.M. solutions $u_{c}$, up to constants, when $\alpha^{\prime}(c)\in\mathbb{R}\setminus\mathbb{Q}$. Thus, we conclude:
Corollary 1.8.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. If $\alpha^{\prime}(c)\in\mathbb{R}\setminus\mathbb{Q}$, then one of the following holds:
  • (1)
    $\mathcal{A}^{*}_{c}$ is a Lipschitz invariant circle.
  • (2)
    $\mathcal{A}^{*}_{c}$ is a Lipschitz Cantori.
1.1. Notations and Hypotheses
@@
  • For the product space $X_{1}\times\cdots\times X_{n}$ and $k\in\{1,\dots,n\}$, the map $\pi_{k}:X_{1}\times\cdots\times X_{n}\rightarrow X_{k}$ is the projection on the $k$-th variable.
@@
  • We denote by $B(\mathbb{T}^{d})$, $C^{0}(\mathbb{T}^{d})$ and $SC(\mathbb{T}^{d})$ the spaces of bounded $\mathbb{Z}^{d}$-periodic functions, continuous $\mathbb{Z}^{d}$-periodic functions and semiconcave $\mathbb{Z}^{d}$-periodic functions.
  • A mapping is of class $C^{1,1}$ if it is differentiable with uniformly Lipschitz derivative.
For the sake of slight generality, we use
@@
    $S$ is ferromagnetic; that is, $S$ is of class $C^{1}$ and the two maps $x\mapsto\partial_{2}S(x,y)$ and $y\mapsto\partial_{1}S(x,y)$ are homeomorphisms between $\mathbb{R}^{d}$ and $(\mathbb{R}^{d})^{*}$ for all $(x,y)$.
Hypothesis 3.
The dimension $d=1$ and the Aubry & Le Daeron Non-crossing Lemma [ALD83] holds: if $(x_{1}-x_{2})(y_{1}-y_{2})<0$, then $S(x_{1},y_{2})+S(x_{2},y_{1})<S(x_{1},y_{1})+S(x_{2},y_{2})$.
@@
One can refer the notion of ferromagnetic terminology in [GT11, Definition 2.4]. If $S$ satisfies Hypothesis 2, then the discrete standard map $F:\mathbb{R}^{d}\times(\mathbb{R}^{d})^{*}\righttoleftarrow$ can be defined by
@@
1.2. Plan for this paper
This paper is organized as follows. In Section 2, we first provide several fundamental facts about discrete weak K.A.M. theory on tori of any dimension.
In Section 2.1, we discuss an important invariant set in discrete weak K.A.M. theory—the Aubry set. In Section 2.2, we address some regularity results in discrete weak K.A.M. theory, particularly those related to semiconcave functions. The conclusions in Section 2.1 require Hypothesis 1, while those in Section 2.2 require both Hypothesis 1 and Hypothesis 2.
Theorems 1.1-1.2 and Theorems 1.4-1.7 are proved in Sections 3 and 4, respectively. The conclusions in Sections 3 and 4 require the validity of Hypothesis 1, Hypothesis 2, and Hypothesis 3.
In Appendix A, we list some fundamental properties of semiconcave functions without proof. Appendix B reviews key results for classical exact twist maps. The proofs of Lemmas 2.19 and 2.22 are included in Appendix C for self-containedness.
2. Discrete weak K.A.M. theory on $d$-dimensional torus
2.1. A notion of important invariant set
As standard, we introduce the following invariant sets and investigate several properties of them. In Section 2.1 , we operate under the assumption that the function $S$ satisfies Hypothesis 1.
We recall the discrete backward Lax-Oleinik operator and discrete forward Lax-Oleinik operator defined in the introduction:
Definition 2.1.
The discrete backward Lax-Oleinik operator $T^{-}$ and discrete forward Lax-Oleinik operator $T^{+}$ are defined by
$\forall y\in\mathbb{R}^{d},\quad T^{-}[u](y):=$ $\inf_{x\in\mathbb{R}^{d}}\big{\{}u(x)+S(x,y)\big{\}},$